3(x+y)- 4 (y-x)= 19
4(2x-y) + y = 18

Answer by stanbon(75887)

3(x+y)- 4 (y-x)= 19
3x+3y -4y + 4x = 19
7x - y = 19
-----------------------
4(2x-y) + y = 18
8x -4y + y = 18
8x-3y = 18
-----------------------
Solve this system:
7x - y = 19
8x - 3y= 18
----------------------
Solve for "y" in the 1st equation:
y = 7x-19
--------------
Substitute for "y" in the 2nd equation and solve for "x":
8x -3*(7x-19) = 18
---
8x - 21x + 57 = 18
-13x = -39
x = 3
-------------------
Since y = 7x-19, so y = 7*3-19 = 2
======================================
Cheers,
Stan H.
===========================